# LLMAID: Identifying AI Capabilities in Android Apps with LLMs

### v1

@@ Abstract.
To address the limitations of existing methods, we propose LLMAID (Large Language Model for AI Discovery), a semantics-driven framework that combines program analysis techniques with large language models (LLMs) to identify and analyze AI services in mobile applications. LLMAID includes four main tasks: (1) candidate extraction, (2) knowledge base interaction, (3) AI capability analysis and detection, and (4) AI service summarization. We apply LLMAID to a dataset of 4,201 Android applications and demonstrate that it identifies 242% more real-world AI apps than state-of-the-art rule-based approaches. Our experiments show that LLM4AID achieves high precision and recall, both exceeding 90%, in detecting AI-related components. Additionally, a user study indicates that developers find the AI service summaries generated by LLMAID to be more informative and preferable to the original app descriptions. Finally, we leverage LLMAID to perform an empirical analysis of AI capabilities across Android apps. The results reveal a strong concentration of AI functionality in computer vision (54.80%), with object detection emerging as the most common task (25.19%).
1. Introduction
@@
Advancements in artificial intelligence (AI) and their successful integration into mobile applications (apps) have enhanced the capabilities of Android apps in daily activities, such as optical character recognition, autonomous driving, object recognition and shopping recommendations. To support developers’ integration of AI capabilities, several vendors have introduced mobile AI frameworks, such as TensorFlow Lite (abadi2016tensorflow), Caffe2 (caffe2:2023), Core ML (coreml:2023), and ncnn (ncnn:2023). These frameworks enable efficient on-device model building and inference, offering better privacy protection and reducing bandwidth usage compared to cloud-based models (guo2018cloud), while also improving responsiveness and ensuring functionality in resource-constrained environments typical of mobile devices. In recent years, there has been a surge in the development of large AI models, such as stable diffusion (stable:2023), ChatGPT (chatgpt:2023), LLaMA (touvron2023llama), and DeepSeek (liu2024deepseek). Unlike traditional lightweight frameworks, these models provide more advanced capabilities in reasoning, multimodal content understanding, and generative tasks. With their advanced abilities to understand and generate content, emerging app developers increasingly resort to these models via the cloud to enhance their businesses (xu2019first; hu2023first), enabling the creation of richer user experiences such as personalized recommendations, interactive assistants, and multimodal content generation.
With the widespread adoption of AI and the growing awareness of its inherent limitations, issues of AI discoverability and transparency have attracted attention from both application consumers and various regulatory bodies. To have a comprehensive understanding of AI services in software artifacts for these stakeholders, this study first tries to uncover the current status quo of AI services in real-world Android applications. In this paper, we define this process as AI Discovery, which encompasses determining whether AI capabilities are integrated or not and further elucidating the functional characteristics of the incorporated AI services. Several studies have been conducted to investigate AI techniques in smartphone apps (xu2019first; zhao2022survey; li2022ai; siu2023towards; deng2019deep). Specifically, Deng et al. (deng2019deep) and Zhao et al. (zhao2022survey) both provided comprehensive surveys on deep learning for mobile devices, highlighting hardware architectures, software frameworks, and optimization mechanisms such as pruning, quantization, and hardware accelerators that enable efficient, privacy-preserving, and cost-effective on-device AI across diverse application domains. Siu et al. (siu2023towards) empirically investigated user–AI interaction in 176 mobile apps, identifying 255 AI features and 759 implementations, ultimately categorizing interaction patterns and demonstrating the importance of usable interaction design for on-device AI features. Li et al. (li2022ai) conducted the largest empirical study to date, analyzing 56,682 AI apps with their newly proposed rule-based approach, analyzing application trends, framework adoption, model protection, and user-related concerns, and releasing an open dataset for future research. Xu et al. (xu2019first) examined 16,500 popular Android apps to reveal early adopters of deep learning, their usage scenarios, and model characteristics, offering insights that emphasize both the potential of mobile deep learning and the need for optimizations, model protection, and validation.
@@
However, these studies exhibit several notable constraints: 1) lacking attention to other potential AI components, [^1] An AI component may refer to either high-level modules (e.g., task planning engines) or low-level implementation artifacts (e.g., APIs, libraries, HTTP requests), whereas an AI feature denotes high-level functionalities (e.g., image recognition, sentiment analysis) enabled by these components. such as integrated development packages, internal assets, and HTTP requests and API invocations in detailed implementations and integrated libraries. Ignoring these components would lead to an incomplete and insufficient analysis; 2) relying on manual inspection or pre-defined keyword searches, which require up-to-date human expertise, making them expensive and difficult to scale; 3) depending on human-defined rule-based methods, which are constrained by domain knowledge and thus ineffective for identifying AI components in diverse real-world applications.
@@
To address these aforementioned challenges, we argue that automating the human-like semantic-based AI Discovery is essential, enabling the characterization of real-world AI components by interpreting their meaningful names. Inspired by recent findings (hagendorff2023human; ning2024can; li2024llms; pan2024large) that large language models (LLMs) are capable of exhibiting human-like reasoning and decision-making, we explore how these models can enhance the AI Discovery process by interpreting semantically meaningful AI component names. Moreover, recent studies have shown that these LLMs can automate various software tasks, including automatic program repair (zhang2022repairing; jiang2023impact; xia2023conversational; han2024chase), code generation (nijkamp2022conversational; chakraborty2022natgen; fried2022incoder; liao20243), and improving software analysis efficiency and effectiveness (ma2023scope; li2023hitchhiker). These findings provide evidence that LLMs may have substantial knowledge in software practice, motivating us to explore their full potential in AI Discovery within mobile applications.
@@
We implemented LLMAID using the widely adopted GPT-3.5-turbo model, which is much more cost-effective than the GPT-4 model (gptmodel:2023). To evaluate the effectiveness of LLMAID, we conducted validation on a sample of 382 AI applications selected from a total of 56,682 already identified AI apps (li2022ai). Manual evaluation of the results confirms that LLMAID achieves high precision and recall (exceeding 90%) in detecting AI components. Furthermore, we applied LLMAID to a dataset of 3,819 real-world Android apps published on Google Play between 2019 and 2023. Compared to the state-of-the-art (SOTA) detection approach, AI Discriminator (li2022ai), LLMAID identifies more than twice as many AI Android apps, showcasing its superior detection capability. Our analysis reveals that the majority of detected AI components pertain to the domains of computer vision (54.80%) and data analysis (26.85%), with object detection (25.19%) and data processing (23.17%) being the most prevalent tasks. At the application level, we find that 47.44% (643/1366) of AI-enabled apps offer services related to data analysis, while 46.41% (634/1366) incorporate computer vision functionalities. We further demonstrate the practical utility of LLMAID by evidencing that developers perceive its AI service report as both informative and preferable.
@@
Figure 1. The workflow of LLMAID.
2. Approach
Fig. 1 illustrates the general framework and workflow of LLMAID (LLM for AI Discovery). As mentioned above, the LLMAID framework consists of four stages: candidate extraction, knowledge base interaction, AI capability analysis and detection, and AI service summarization. Together, these stages are designed to address the complexity of uncovering AI functionalities in Android apps. In the first stage, Candidate extraction, static analysis techniques are applied to the Android APK to extract fine-grained components that may potentially involve AI functionality. The second stage, Knowledge base interaction, refines these candidates by referencing dynamically constructed AI and non-AI knowledge bases, which act as caches of previously analyzed components. This step reduces redundant analysis and improves precision by filtering out irrelevant components early. The third stage, AI capability analysis and detection, prompts a large language model to interpret the functionality of each remaining candidate and determine whether it constitutes an AI component, yielding a curated set of AI-relevant elements together with functional descriptions. Finally, the AI service summarization stage produces a concise AI service report for the given APK by leveraging a carefully designed LLM prompt, which incorporates the determined AI components along with their interpreted AI functional descriptions. Importantly, the overall workflow is model-agnostic and can be applied with any advanced large language model, ensuring adaptability as LLM technologies continue to evolve. We elaborate on each stage in the following subsections.
Figure 2. The structure of a typical released Android APK.
2.1. Task 1: Candidate Extraction
To determine AI capabilities within an Android application, we first extract all relevant information—referred to as candidates-from the APK file. Fig. 2 shows the structure of a typical released Android APK file, typically including classes.dex, which is the Dalvik executable file (dalvik:2022), assets, META-INFO, lib, AndroidManifest.xml, res, and resources.arsc. Notably, an APK may comprise multiple .dex files because each file is constrained by the 65,536-method limit. When this threshold is exceeded, the compiler automatically partitions the code into additional files, conventionally named classes1.dex, classes2.dex, and so forth.
@@ 2.1. Task 1: Candidate Extraction
We utilize AAPT (aapt:2024) to extract the structure from the curated Android APK file. Our focus is on examining the implementation details embedded in the app, which requires decompiling .dex files to retrieve the underlying Java logic and analyzing .so files in the lib directory for native code. Specifically, we extend a static analysis approach, namely App Scanner, proposed by Liu et al. (liu2021identifying) to extract APIs from an APK file. The scanner leverages the Soot framework (lam2011soot), a widely used Java-based static analysis framework that has been extended to support Android applications. In particular, Soot translates Dalvik bytecode (dalvik:2022), the default executable format for Android applications generated from Java (or Kotlin) source code through an intermediate compilation step and optimized for devices with limited memory and power resources, into Jimple (bartel2012dexpler). Jimple is a three-address intermediate representation that is both human-readable and well suited for static program analysis. This transformation bridges the gap between low-level Dalvik instructions and a higher-level abstraction and enables a more effective analysis of Android applications. The App Scanner method in this study is enhanced to additionally capture fine-grained components from multiple sources within the given APK, such as .dex files, native library .so files, and the asset files. This also allows us to extract more indicators of AI functionality, such as HTTP requests https://api.openai.com/v1/ and descriptive package name com.google.firebase.ml.vision.objects. Moreover, to analyze embedded dynamic native libraries, we employ tools readelf (readelf:2023) and objdump (objdump:2023) to parse HTTPS requests, as many apps implement AI capabilities by invoking these requests via the Android NDK (ndk:2023).
2.2. Task 2: Knowledge Base Interaction
Given our focus on implementation-level details of AI capabilities, the extraction process produces a substantial number of fine-grained components that include both AI-relevant and non-AI functionalities. The presence of non-AI components introduces noise into the analysis, as they do not contribute to the understanding of AI capabilities and can impede the accurate identification of AI-related functionality. To reduce such noise, we first apply a rule-based filtering approach designed to exclude non-AI components, such as official Java and Android kernel-related; however, because this method operates at a coarse granularity, some non-AI components may still be incorrectly retained as AI-related.
To address the limitations of coarse-grained rule-based filtering and further improve analysis efficiency, we introduce a knowledge base (KB) interaction step into the workflow. This step enables the reuse of analysis outcomes for components that have already been processed, allowing previously verified AI or non-AI components to be retrieved directly and thereby reducing redundant computation. To support this process, we construct two complementary KBs that function as caches rather than shortcuts for classification: an AI KB, which stores verified AI components together with their capability analyses, and a non-AI KB, which records confirmed non-AI components to facilitate pre-analysis filtering. After component extraction and initial filtering, the system checks each candidate against both KBs; if a match is found, the corresponding analysis and detection result can be directly retrieved, thereby bypassing further inference. Otherwise, the components are forwarded to the next stage, Task 3: AI capability analysis and detection, whose results are then systematically incorporated into the corresponding KB for future reference and reuse. To ensure reliability, both KBs are continuously refined and manually audited, ensuring that cached results maintain high accuracy while significantly reducing unnecessary computation.
2.3. Task 3: AI Capability Analysis and Detection
@@
This stage, the core of LLMAID, focuses on identifying which of the extracted components are AI-related and characterizing them in terms of their AI functional capabilities. To ensure precision and interpretability, this task is divided into two sub-steps: AI capability analysis and AI component determination. The first sub-step, AI capability analysis, interprets the functional behavior of each candidate component to identify the type of functionality it provides, such as image recognition, language processing, or network communication. The second sub-step, AI component determination, then decides whether the interpreted component should be classified as AI-related, filtering out non-AI elements. After the knowledge base interaction phase, we obtain a set of deduplicated candidate components that serve as the input for this two-step process. For each candidate, we issue an initial prompt to ChatGPT to generate a functional-oriented interpretation that supports both capability analysis and component determination. For example, given the package name com.google.mlkit.vision.objects, ChatGPT may provide the analysis result of “Google’s ML Kit API for object detection and tracking in images and videos.” In cases where components are obfuscated or highly customized, i.e., nlp.WordPieceModelPB, ChatGPT can still offer meaningful functional descriptions, e.g., “A library for tokenizing text using the WordPiece algorithm, implemented in Protocol Buffers format” which would be difficult to identify using a rule-based approach. Based on the generated analysis, we further prompt ChatGPT to determine whether the component falls within the scope of AI and justify its decision with a detailed explanation. Leveraging ChatGPT’s ability to interpret and reason over its own generated content, the sub-steps of AI capability analysis and AI component determination can be straightforwardly and effectively conducted by adopting the prompts of Task 3.1 and Task 3.2 in Fig. 3.
2.4. Task 4: AI Service Summarization
Figure 3. Prompt design and examples for LLMAID
@@ 2.4. Task 4: AI Service Summarization
Given the determined AI components together with their corresponding functional descriptions, we construct and employ a carefully designed ChatGPT prompt to generate a concise AI service report. This report synthesizes the functional interpretations of the identified components into a coherent summary of their AI capabilities, thereby providing an interpretable overview that bridges low-level implementation details with higher-level functional insights. Nevertheless, due to the inherent context-length limitations of ChatGPT, large language models (LLMs) in general, it is often infeasible to incorporate all component names and their corresponding analyses into a single prompt, which necessitates additional strategies to manage input complexity. To mitigate this constraint, we exclude low-level API components whose functionalities can generally be inferred from their package context, and instead concentrate on higher-level components. The remaining AI-relevant components and their interpreted AI functional descriptions are concatenated and incorporated into a carefully constructed prompt for ChatGPT, which is then used to generate the AI service report, as illustrated in Task 4 of Fig. 3. It is worth noting that the generated report can be tailored to different audiences, such as end users, regulators, or developers, by making minor adjustments to the prompting strategy, thereby enabling context-specific interpretations and supporting diverse analytical needs.
2.5. Experiment Configuration
@@
In this study, to instruct ChatGPT to execute each sub-step precisely, we carefully design natural language prompts, which are presented in Fig. 3. To further guide ChatGPT in understanding the intended analysis tasks, we provide five input–output examples to enable few-shot in-context learning(brown2020language; wang2020generalizing). In our implementation of LLMAID, we utilize the default 4K-token context window provided by the ChatGPT API. However, due to ChatGPT’s rate limit, processing one app with one component per request is both time-consuming and inefficient. Conversely, submitting too many components in a single request risks output truncation or omission, which may cause misalignment between inputs and outputs. To balance task accuracy and processing efficiency, we adopt a batch size of three components per request and instruct ChatGPT to return structured outputs in JSON format. The model parameters are left at default values, except for the temperature and TopP settings, which are adjusted from 1.0 to 0.2 and 0.95, respectively. These adjustments reduce output variability while retaining a certain degree of creativity, thereby improving the consistency and reliability of the component analyses and the resulting AI service reports.
3. Evaluation

### v2

@@ Abstract
To address the limitations of existing methods, we propose LLMAID (Large Language Model for AI Discovery), a semantics-driven framework that combines program analysis techniques with large language models (LLMs) to identify and analyze AI services in mobile applications. LLMAID includes four main tasks: (1) candidate extraction, (2) knowledge base interaction, (3) AI capability analysis and detection, and (4) AI service summarization. We apply LLMAID to a dataset of 4,201 Android applications and demonstrate that it identifies 242% more real-world AI apps than state-of-the-art rule-based approaches. Our experiments show that LLMAID achieves high precision and recall, both exceeding 90%, in detecting AI-related components. Additionally, a user study indicates that developers find the AI service summaries generated by LLMAID to be more informative and preferable to the original app descriptions. Finally, we leverage LLMAID to perform an empirical analysis of AI capabilities across Android apps. The results reveal a strong concentration of AI functionality in computer vision (54.80%), with object detection emerging as the most common task (25.19%).
1 Introduction
@@
Advancements in artificial intelligence (AI) and their successful integration into mobile applications (apps) have enhanced the capabilities of Android apps in daily activities, such as optical character recognition, autonomous driving, object recognition and shopping recommendations. To support developers’ integration of AI capabilities, several vendors have introduced mobile AI frameworks, such as TensorFlow Lite abadi2016tensorflow, Caffe2 caffe2:2023, Core ML coreml:2023, and ncnn ncnn:2023. These frameworks enable efficient on-device model building and inference, offering better privacy protection and reducing bandwidth usage compared to cloud-based models guo2018cloud. In recent years, there has been a surge in the development of large AI models, such as stable diffusion stable:2023, ChatGPT chatgpt:2023, LLaMA touvron2023llama, and DeepSeek liu2024deepseek. With their advanced abilities to understand and generate content, emerging app developers increasingly resort to these models via the cloud to enhance their businesses xu2019first; hu2023first.
@@
With the widespread adoption of AI and the growing awareness of its inherent limitations, issues of AI discoverability and transparency have attracted attention from both application consumers and various regulatory bodies. To have a comprehensive understanding of AI services in software artifacts for these stakeholders, this study first tries to uncover the current status quo of AI services in real-world Android applications. In this paper, we define this process as AI Discovery, which encompasses determining whether AI capabilities are integrated or not and further elucidating the functional characteristics of the incorporated AI services. Several studies have been conducted to investigate AI techniques in smartphone apps xu2019first; zhao2022survey; li2022ai; siu2023towards; deng2019deep. However, these studies exhibit several notable constraints: 1) lacking attention to other potential AI components, [^1] Component in the paper may refer to either high-level modules or low-level elements such as APIs and libraries, depending on context. such as integrated development packages, internal assets, and HTTP requests and API invocations in detailed implementations and integrated libraries. Ignoring these components would lead to an incomplete and insufficient analysis; 2) relying on manual inspection or pre-defined keyword searches, which require up-to-date human expertise, making them expensive and difficult to scale; 3) depending on human-defined rule-based methods, which are constrained by domain knowledge and thus ineffective for identifying AI components in diverse real-world applications.
@@
To address these aforementioned challenges, we argue that automating the human-like semantic-based AI Discovery is essential, enabling the characterization of real-world AI components by interpreting their meaningful names. Inspired by recent findings hagendorff2023human; ning2024can; li2024llms; pan2024large that large language models (LLMs) are capable of exhibiting human-like reasoning and decision-making, we explore how these models can enhance the AI Discovery process by interpreting semantically meaningful AI component names. Moreover, recent studies have shown that these LLMs can automate various software tasks, including automatic program repair zhang2022repairing; jiang2023impact; xia2023conversational; han2024chase, code generation nijkamp2022conversational; chakraborty2022natgen; fried2022incoder; liao20243, and improving software analysis efficiency and effectiveness ma2023scope; li2023hitchhiker. These findings provide evidence that LLMs may have substantial knowledge in software practice, motivating us to explore their full potential in AI Discovery within mobile applications.
@@
We implemented LLMAID using the widely adopted GPT-3.5-turbo model, which is much more cost-effective than the advanced GPT-4 model gptmodel:2023. To evaluate the effectiveness of LLMAID, we conducted validation on a sample of 382 AI applications selected from a total of 56,682 already identified AI apps li2022ai. Manual evaluation of the results confirms that LLMAID achieves high precision and recall (exceeding 90%) in detecting AI components. Furthermore, we applied LLMAID to a dataset of 3,819 real-world Android apps published on Google Play between 2019 and 2023. Compared to the state-of-the-art (SOTA) detection approach, AI Discriminator li2022ai, LLMAID identifies more than twice as many AI Android apps, showcasing its superior detection capability. Our analysis reveals that the majority of detected AI components pertain to the domains of computer vision (54.80%) and data analysis (26.85%), with object detection (25.19%) and data processing (23.17%) being the most prevalent tasks. At the application level, we find that 47.44% (643/1366) of AI-enabled apps offer services related to data analysis, while 46.41% (634/1366) incorporate computer vision functionalities. We further demonstrate the practical utility of LLMAID by evidencing that developers perceive its AI service report as both informative and preferable.
@@
Figure 1: The workflow of LLMAID.
2 Approach
Fig. 1 illustrates the general framework and workflow of LLMAID (LLM for AI Discovery). As mentioned above, the LLMAID framework consists of four stages. These stages are designed to address the complexity of uncovering AI functionality in Android apps. Candidate extraction employs static analysis techniques to extract potential AI components from an Android APK. Knowledge base interaction aims to filter these extracted components using a constructed AI and non-AI knowledge base, thereby improving the accuracy of downstream AI capability analysis. The AI capability analysis and detection stage further identifies and characterizes AI components by prompting an LLM to analyze. Finally, AI service summarization automatically generates a report of these AI components for the given Android APK. Importantly, our approach is model-agnostic and can be applied to any advanced large language model. We describe each stage in the following subsections.
Figure 2: The structure of a typical released Android APK.
2.1 Task 1: Candidate Extraction
@@
To determine AI capabilities within an Android application, we first extract all relevant information—referred to as candidates—from the APK file. Fig. 2 shows the structure of a typical released Android APK file, typically including classes.dex, which is the Dalvik executable file dalvik:2022, assets, META-INFO, lib, AndroidManifest.xml, res, and resources.arsc. Notably, an APK may contain multiple .dex files due to the method limit of 65,536 per file. We utilize AAPT aapt:2024 to extract the structure from the curated Android APK file. Our focus is on examining the implementation details embedded in the app, which requires decompiling .dex files to retrieve the underlying Java logic and analyzing .so files in the lib directory for native code. Specifically, we extend a static analysis approach, namely App Scanner, proposed by Liu et al. liu2021identifying to extract APIs from an APK file. The scanner aims to transform Dalvik bytecode dalvik:2022 into Jimple bartel2012dexpler, a human-readable intermediate representation. The App Scanner method is enhanced to additionally capture fine-grained components from multiple sources within the given APK, such as .dex files, native library .so files, and the asset files. This also allows us to extract more indicators of AI functionality, such as HTTP requests https://api.openai.com/v1/ and descriptive package name com.google.firebase.ml.vision.o-
bjects
. Moreover, to analyze embedded dynamic native libraries, we employ tools readelf readelf:2023 and objdump objdump:2023 to parse HTTPS requests, as many apps implement AI capabilities by invoking these requests via the Android NDK ndk:2023.
Figure 3: Prompt design and examples for LLMAID
2.2 Task 2: Knowledge Base Interaction
The extraction process yields a substantial number of fine-grained components capturing both AI and non-AI implementation details. To filter out the irrelevant non-AI components, we apply a rule-based approach following the extraction process. This approach aims to remove components that are irrelevant to AI functionalities; however, it operates at a coarse granularity, which may result in the retention of some non-AI components falsely classified as AI-related. To further refine the analysis and enhance efficiency, we incorporate a knowledge base (KB) interaction step into the workflow. This step enables the reuse of prior analysis results to accelerate processing and reduce redundant computation. Specifically, we maintain two separate KBs: an AI KB that stores previously verified AI components and their associated capability analyses, and a non-AI KB that contains confirmed non-AI components used primarily for pre-analysis filtering. Once the components are extracted and initially filtered, they are checked against both KBs. If a match is found, the corresponding analysis and detection result are retrieved directly, bypassing further computation. Otherwise, the components are forwarded to the next stage, Task 3: AI capability analysis and detection, whose results are then systematically incorporated into the corresponding KB for future reference and reuse.
2.3 Task 3: AI Capability Analysis and Detection
@@
This stage, the core of LLMAID, involves identifying and characterizing the extracted AI components. To ensure precision and interpretability, this task is divided into two sub-steps: AI capability analysis and AI component determination. After passing through the knowledge base interaction phase, we obtain a set of deduplicated candidate components that are used for further analysis. For each component, we issue an initial prompt to ChatGPT to produce a functional-oriented interpretation. For example, given the package name com.google.mlkit.vision.objects, ChatGPT may provide the analysis result of “Google’s ML Kit API for object detection and tracking in images and videos.” In cases where components are obfuscated or highly customized, i.e., nlp.WordPieceModelPB, ChatGPT can still offer meaningful functional descriptions, e.g., “A library for tokenizing text using the WordPiece algorithm, implemented in Protocol Buffers format” which would be difficult to identify using a rule-based method approach. Based on the generated analysis, we further prompt ChatGPT to determine whether the component falls within the scope of AI and justify its decision with a detailed explanation. Leveraging ChatGPT’s ability to interpret and reason over its own generated content, the sub-steps of AI capability analysis and AI component determination can be straightforwardly and effectively conducted by adopting the prompts of Task 3.1 and Task 3.2 in Fig. 3.
2.4 Task 4: AI Service Summarization
@@
Given the determined AI components and their corresponding functional descriptions, we generate a concise AI service report by employing a a well-crafted ChatGPT prompt to summarize the AI capabilities of these components. However, due to the inherent context length limitations of ChatGPT and LLMs in general, incorporating all component names and their analyses into a single prompt is often infeasible. To mitigate this constraint, we omit low-level API components, i.e., their functionalities are typically inferred from their package context, and focus on higher-level components instead. By concatenating the remaining AI-relevant components, we prompt ChatGPT to generate the AI service report for app users, which is shown in Task 4 of Fig. 3. Notably, the generated report can be adapted for different audiences, such as end users, regulators, or developers, by slightly adjusting the prompt, thereby enabling context-specific interpretations and supporting diverse analysis needs.
2.5 Experiment Configuration
@@
To instruct ChatGPT to execute each sub-step precisely, we carefully design natural language prompts, which are presented in Fig. 3. To make ChatGPT fully understand these analysis steps, we additionally provide five input-output examples for the few-shot in-context learning brown2020language; wang2020generalizing. In our implementation of LLMAID, we utilize the default 4K-token context window provided by the ChatGPT API. Given ChatGPT’s rate-limiting, processing one app with one component per request becomes time-consuming and inefficient. Moreover, submitting too many components in a single request may lead to output truncation or omission, resulting in misalignment between input and output. To balance task accuracy and processing efficiency, we adopt a batch size of three components per request and instruct ChatGPT to return structured outputs in JSON format. Model parameters are left at default values, except for the temperature and TopP settings, which are adjusted from 1.0 to 0.2 and 0.95, respectively. These adjustments reduce output variability while retaining a certain degree of creativity, thereby improving the consistency and reliability of the component analyses and summary reports.
3 Evaluation